The mean weight of loads of rock is 47.0 tons with a standard deviation of 6.0 tons. If 16 loads are chosen at random for a weight check, find the probability that the mean weight of those loads is less than 46.2 tons. Assume that the variable is normally distributed.




1
Expert's answer
2021-06-22T14:19:27-0400

"\mu = 47.0 \\

\sigma=6.0 \\

n=16 \\

P(\bar{X}<46.2) = P(Z< \frac{46.2-47}{6.0/ \sqrt{16}} )\\

= P(Z< -0.533) \\

= 0.2970"